OpenStudy (anonymous):

Solve the system by the elimination method. 2x + y - 4 = 0 2x - y - 4 = 0 When you eliminate y, what is the resulting equation? x = 0 4x = -8 4x = 8

jimthompson5910 (jim_thompson5910):

What do you get when you add up the equations 2x + y - 4 = 0 2x - y - 4 = 0

OpenStudy (anonymous):

can you explain?

jimthompson5910 (jim_thompson5910):

notice how +y added to -y gives you y + (-y) = y - y = 0y = 0

jimthompson5910 (jim_thompson5910):

so that means when you add the equations up, the y terms will cancel out

jimthompson5910 (jim_thompson5910):

so what happens when you add up the x terms?

OpenStudy (jdoe0001):

I assuming you are wanting to understand it, rather than the "exercise asks you to explain it"

OpenStudy (jdoe0001):

because it was rather clear in the previous posting